# XtraLight-MedMamba for Classification of Neoplastic Tubular Adenomas

### v3

@@ Abstract
Accurate risk stratification of precancerous polyps during routine colonoscopy screenings is essential for lowering the risk of developing colorectal cancer (CRC). However, assessment of low grade dysplasia remains limited by subjective histopathologic interpretation. Advancements in digital pathology and deep learning provide new opportunities to identify subtle and fine morphologic patterns associated with malignant progression that may be imperceptible to the human eye. In this work, we propose XtraLight-MedMamba, an ultra-lightweight state-space–based deep learning framework for classifying neoplastic tubular adenomas from whole-slide images (WSIs). The architecture is a blend of ConvNext based shallow feature extractor with parallel vision mamba to efficiently model both long- and short-range dependencies and image generalization. An integration of Spatial and Channel Attention Bridge (SCAB) module enhances multiscale feature extraction, while Fixed Non-Negative Orthogonal Classifier (FNOClassifier) enables substantial parameter reduction and improved generalization. The model was evaluated on a curated dataset acquired from patients with low-grade tubular adenomas, stratified into case and control cohorts based on subsequent CRC development. XtraLight-MedMamba achieved an accuracy of 97.18% and an F1-score of 0.9767 using approximately 32,000 parameters, outperforming transformer-based and conventional Mamba architectures with significantly higher model complexity.
@@
Digital Pathology, tubular adenoma, Colorectal Cancer, parallel vision mamba, state space models, ConvNext, lightweight deep learning, whole slide images
@@ 1 Introduction
Colorectal cancer (CRC) remains a major global health challenge. It is the third most commonly diagnosed cancer and the second leading cause of cancer-related mortality in the United States [22],[21]. Colonic polyps are raised protrusions of colonic mucosa, and different types have the potential to progress to cancer. CRC most commonly arises from adenomatous polyps, well-characterized precursor (premalignant) lesions of the colonic mucosa resulting from neoplastic proliferation of glandular colonic epithelium. Although these types of polyps are benign, they carry malignant potential, capable of progressing to cancer via the adenoma-carcinoma sequence in a stepwise progression [18],[1],[28]. Over decades, the classical adenoma–carcinoma sequence has established adenomatous polyps as a central biological intermediate in colorectal tumorigenesis, forming the basis for modern CRC screening and surveillance strategies [28],[4].
@@
Figure 1: Schematic illustrating the concept of the adenoma–carcinoma sequence described by Fearon and Vogelstein [28]. Simply put, the central idea is that colorectal cancer develops in a step-wise progressive manner, starting from the precursor lesions (adenomatous polyps), eventually progressing to invasive carcinoma. Early loss of APC is associated with hyperproliferative epithelium, increasing the risk of colonic polyp formation. This is followed by an activating mutation in KRAS, which itself promotes adenomatous polyp(s) formation. Subsequent loss of tumor suppressor genes, including TP53 and DCC, contributes to tumor progression and eventual neoplastic progression towards invasive colorectal adenocarcinoma [28]. This step-wise progression from adenoma(s) to cancer typically takes over ten years to progress [15].
@@
Amongst the adenomatous polyps are subtypes. Tubular adenomas (TAs) represent the most prevalent subtype encountered in routine clinical practice and are therefore of particular clinical importance. The adenomatous polyps are histologically defined by their degree of dysplasia (abnormal tissue architecture) into low-grade and high-grade. In general, high-grade dysplasia is a characteristic predictor of advancement to colorectal carcinoma [18]. The majority of TAs detected via screening colonoscopy and histopathological examination typically demonstrate low-grade dysplasia, in which malignant potential is low-risk [18],[15]. Current histopathologic assessment relies primarily on subjective visual interpretation, which may be insufficient to identify subtle or spatially distributed morphologic features associated with long-term development of CRC in this ostensibly low-risk population. Epidemiologic trends further highlight the importance of improving risk stratification at the precursor lesion stage [22],[1]. Despite widespread screening initiatives among adults aged 45+ years reducing the overall incidence of CRC, important age-related disparities continue to exist; in recent years, incidence has continued to rise among younger individuals despite overall declines [22],[4]. Historical and recent data continues to highlight the continued clinical relevance of early detection and limitations in current risk assessment strategies based on coarse histologic categorization. Despite the success of screening programs in reducing CRC incidence and mortality, a significant unmet clinical need remains in accurately stratifying risk among patients with low-grade tubular adenomas. Most individuals diagnosed with TAs are classified as low risk and managed conservatively, yet a subset will continue to later develop advanced neoplasia or CRC. Continued progress in CRC prevention will therefore require improved tools capable of extracting objective, quantitative information from precursor lesions beyond what is achievable through conventional histopathologic evaluation alone. The increasing adoption of digital pathology has transformed histopathology into a data-rich imaging modality through the routine acquisition of high-resolution whole-slide images (WSIs). When coupled with advances in machine learning, digital pathology offers the potential to enhance prophylactic efforts by enabling large-scale, quantitative analysis of tissue morphology and by uncovering subtle patterns imperceptible to human observers [10],[13]. However, the high resolution, heterogeneity, and weak labeling of WSI data introduce substantial computational challenges, necessitating efficient, scalable analytical frameworks.
@@
Recent advances in computational pathology and deep learning architectures have enabled the automated depiction of features in the WSIs, allowing models to learn the hierarchical representation of tissue patterns directly from histolgic data without the need for manual intervention. With early efforts driven by convolutional neural network (CNN)-based models [14], [he_deep_2016], [9], [27] and subsequently transformer-inspired models [11], [3] strong performance has been achieved by modeling local textural features and global context through self-attention mechanisms, respectively. While effective, CNNs have limited receptive field constraining their ability to model long-range spatial relationships that are critical for understanding broader tissue architecture. Whereas, transformers incur computational overhead due to quadratic complexity, posing challenges for WSI analysis tasks.
@@
Most recently, state space models (SSMs), have introduced new opportunities for efficient representation learning in large-scale visual data [6, 5]. Vision Mamba, [31, 17, 26] an SSM-based architecture, enables modeling of short- and long-range spatial dependencies with linear computational complexity, all while maintaining favorable computational and memory characteristics, making it well-suited for WSI analysis. In this context, the application of modern, computationally efficient architectures to digital pathology represents a critical step toward advancing cancer prevention through improved risk stratification of precursor lesions.
@@
Conventional risk stratification methods include the categorization of adenomas based on size, histologic subtype, and grade of dysplasia, potentially overlooking finer-scale tissue patterns that may reflect underlying biological behavior. Tubular adenomas with low-grade dysplasia are often regarded as biologically homogeneous, despite evidence that not all precancerous lesions progress to malignant disease. Subtle morphologic cues and variations, such as nuclear architecture, glandular organization, and epithelial maturation, may be difficult for human observers to consistently quantify but can be captured by deep learning models. This paper proposes XtraLight-MedMamba, an ultra-lightweight parallel Mamba-based network for the classification of neoplastic tubular adenomas into “case” and “control” cohorts by enhancing and identifying subtle morphologic pattern cues through the modeling of both short- and long-range spatial dependencies [25]. Through testing and evaluation, the new model, XtraLight-MedMamba, demonstrates superior performance compared to previously established models, as detailed in subsequent sections.
@@
  • We propose a novel mamba-based extra light architecture for classification neoplastic tubular adenomas: XtraLight-MedMamba, combining the strengths of convNext block, mamba modules in parallel layers, spatial and channel attention bridge (SCAB) module and fixed non-negative orthogonal classifier (FNOClassifier)
@@
  • The novel architecture uses (1) ConvNext blocks as a shallow feature extractor, acting upon design principles of Vision Transformers [dosovitskiy_image_2021] (ViT) such as, depthwise separable convolutions, Layer Normalization, and MLP-style channel expansion to efficiently capture local morphologic patterns while preserving the spatial inductive bias of CNNs. (2) Parallel Vision Mamba (PVM) layers are included to parallely model both short- and long-range spatial dependencies by branching out the features for state-space sequence modeling each with channel $C/4$, enabling efficient global contextualization of morphologic features with linear computational complexity. (3) Multiscale feature fusion is enhanced by integration of SCAB modules by emphasizing relevant informative spatial regions and channel-wise responses. (4)
  • FNOClassifier is incorporated to reduce redundant parameters deeper in the layers of the model while improving generalization and training stability through structured, orthogonal feature-to-class projections.
  • The effectiveness of the proposed XtraLight-MedMamba model is validated using the curated WSIs, tiled RGB images from Neoplastic Tubular Adenoma (NPTA) dataset [25].
@@
By focusing on morphologic features critical for tissue characterization, XtraLight-MedMamba is aimed at improving the accuracy of neoplastic tubular adenoma classification by capturing subtle cellular and complex architectural patterns present in whole-slide images from case and control cohorts.
2 Related Works
Computational analysis of histopathology images has become a key research focus for improving CRC diagnosis and evaluation, driven by advances in computer vision. Early computational approaches relied on manually-developed algorithms to capture image features like texture [hamilton_automated_1997] or image histograms [kalkan_automated_2012, sengar_grading_2016], followed by classical supervised learning methods like support vector machines [kather_multi-class_2016, sengar_grading_2016] or k-nearest neighbors [kalkan_automated_2012]. While these methods demonstrated the feasibility of automated CRC classification, they depended on human-selected features which may not generalize well [hamilton_automated_1997] and their performance was generally lower than human pathologists [kalkan_automated_2012, sengar_grading_2016].
The emergence of CNNs for computer vision, following the success of AlexNet in 2012 [krizhevsky_imagenet_2012], spurred their widespread adoption in digital pathology, including for CRC pathologies [ben_hamida_deep_2021, davri_deep_2022]. CNNs have been employed for tasks like cancer detection [song_automatic_2020, wang_accurate_2021, haq_yolo_2024], prognosis prediction [zhou_construction_2024, trahearn_computational_2025], and tumor grading [24, kim_detection_2023, perlo_dysplasia_2022], with performances competitive with human pathologists. More recently, Transformer-based architectures, enabled by the introduction of the Vision Transformer [dosovitskiy_image_2021], have been explored to better capture long-range spatial relationships within colorectal tissue [sathyanarayana_colovit_2025, tp_deepcpd_2024, qin_colorectal_2024], and for detection of higher-order structures within CRC [le_rochais_automated_2025, chen_predicting_2024]. However, transformer-based architectures rely heavily on the self-attention mechanism to capture these subtle patterns, which exhibits quadratic growth [6, keles_computational_2023], thus making them less suitable for use on gigapixel-sized WSIs.
Subsequently, SSM-based architectures [6] like Vision Mamba [31] have been used to great effect in digital pathology tasks like WSI image classification [zhang_2dmamba_2025], due to their ability to capture global and local spatial context with far less computational power than transformers [zheng_m3amba_2025, ding_combining_2025]. However, most research using SSMs for histopathology tasks has been in more well-studied pathologies like lung cancer [khan_slidemamba_2026, ding_combining_2025] and breast cancer [malekmohammadi_classification_2024, nasiri-sarvi_vim4path_2024], rather than CRC. As far as we are aware, this is the first work using a mamba-based architecture for neoplastic tubular adenoma classification in CRC histopathology slides.

### v4

@@ Abstract
Accurate risk stratification of precancerous polyps during routine colonoscopy screening is a key strategy to reduce the incidence of colorectal cancer (CRC). However, assessment of low-grade dysplasia remains limited by subjective histopathologic interpretation. Advances in computational pathology and deep learning offer new opportunities to identify subtle, fine morphologic patterns associated with malignant progression that may be imperceptible to the human eye. In this work, we propose XtraLight-MedMamba, an ultra-lightweight state-space–based deep learning framework to classify neoplastic tubular adenomas from whole-slide images (WSIs). The architecture is a blend of a ConvNeXt-based shallow feature extractor with parallel vision mamba blocks to efficiently model local texture cues within global contextual structure. An integration of the Spatial and Channel Attention Bridge (SCAB) module enhances multiscale feature extraction, while the Fixed Non-Negative Orthogonal Classifier (FNOClassifier) enables substantial parameter reduction and improved generalization. The model was evaluated on a curated dataset acquired from patients with low-grade tubular adenomas, stratified into case and control cohorts based on subsequent CRC development. XtraLight-MedMamba achieved an accuracy of 97.18% and an F1-score of 0.9767 using approximately 32,000 parameters, outperforming transformer-based and conventional Mamba architectures, which have significantly higher model complexity and computational burden, making it suitable for resource-constrained areas.
@@
Digital pathology, tubular adenoma, Colorectal Cancer, parallel vision mamba, state space models, ConvNeXt, lightweight deep learning, whole-slide images
@@ 1 Introduction
Colorectal cancer (CRC) remains a major public-health burden worldwide. It is the third most commonly diagnosed cancer and the second leading cause of cancer-related mortality in the United States [37],[36]. Colonic polyps are raised protrusions of colonic mucosa, comprising several histologic subtypes with differing biological behavior. Among these, adenomatous polyps represent a major category of premalignant lesions arising from the neoplastic proliferation of colonic glands. Neoplasia refers to abnormal clonal cell growth that persists in the absence of normal regulatory signals. Although these types of polyps are typically benign, this autonomous proliferative capacity is why adenomatous polyps are considered precancerous lesions and can progress to cancer via the adenoma-carcinoma sequence in a stepwise manner [32],[2],[44] as shown in Fig. 1. Over decades, the classical adenoma–carcinoma sequence has established adenomatous polyps as a central biological intermediate in colorectal tumorigenesis, forming the basis for modern CRC screening and surveillance strategies [44],[9].
@@
Figure 1: Schematic illustrating the concept of the adenoma–carcinoma sequence described by Fearon and Vogelstein [44]. Together, these observations suggest that the central idea is that colorectal cancer develops in a step-wise progressive manner, starting from the precursor lesions (adenomatous polyps), eventually progressing to invasive carcinoma. Early loss of APC is associated with hyperproliferative epithelium, increasing the risk of colonic polyp formation. This is followed by an activating mutation in KRAS, which itself promotes the formation of adenomatous polyps. Subsequent loss of tumor suppressor genes, including TP53 and DCC, contributes to tumor progression and eventual neoplastic progression towards invasive colorectal adenocarcinoma [44]. This step-wise progression from adenoma(s) to cancer typically takes over ten years to progress [26].
@@
Amongst the adenomatous polyps are subtypes. Tubular adenomas (TAs) represent the most prevalent subtype encountered in routine clinical practice and are therefore of particular clinical importance. Adenomatous polyps are histologically classified as low-grade or high-grade based on the degree of dysplasia (abnormal tissue architecture). In general, high-grade dysplasia is a characteristic predictor of advancement to colorectal carcinoma [32]. The majority of TAs detected via screening colonoscopy and histopathological examination typically demonstrate low-grade dysplasia, in which malignant potential is low-risk [32],[26]. Current histopathologic assessment relies primarily on subjective visual interpretation, which may be insufficient to identify subtle or spatially distributed morphologic features associated with long-term development of CRC in this ostensibly low-risk population. Epidemiologic trends further highlight the importance of improving risk stratification at the precursor lesion stage [37],[2]. Despite widespread screening initiatives among adults aged 45+ that have reduced the overall incidence of CRC, important age-related disparities persist; in recent years, incidence has continued to rise among younger individuals despite overall declines [37],[9]. Historical and recent data continue to highlight the clinical relevance of early detection and the limitations of current risk assessment strategies based on coarse histologic categorization. Despite the success of screening programs in reducing CRC incidence and mortality, a significant unmet clinical need remains in accurately stratifying risk among patients with low-grade tubular adenomas (TAs). Most individuals diagnosed with TAs are classified as low risk and managed conservatively, yet a subset will continue to later develop high-grade dysplasia, neoplasia, and/or CRC. Continued progress in CRC prevention will therefore require improved tools capable of extracting objective, quantitative information from precursor lesions beyond what is achievable through conventional histopathologic evaluation alone. The increasing adoption of digital pathology has transformed histopathology into a data-rich imaging modality through the routine acquisition of high-resolution whole-slide images (WSIs). When coupled with advances in machine learning, digital pathology offers the potential to enhance prophylactic efforts by enabling large-scale, quantitative analysis of tissue morphology and by uncovering subtle patterns imperceptible to human observers [20],[24]. However, the high resolution, heterogeneity, and weak labeling of WSI data introduce substantial computational challenges, necessitating efficient, scalable analytical frameworks.
@@
Recent advances in computational pathology and deep learning architectures have enabled automated feature extraction from WSIs, allowing models to learn hierarchical representations of tissue patterns directly from histological data without manual intervention. With early efforts driven by convolutional neural network (CNN)-based models [25], [he_deep_2016], [19], [42], and subsequently transformer-inspired models [21], [8], strong performance has been achieved by modeling local textural features and global context through self-attention mechanisms, respectively. While effective, CNNs have a limited receptive field, constraining their ability to model long-range spatial relationships that are critical for understanding broader tissue architecture. Whereas transformers incur computational overhead due to their quadratic complexity, which poses challenges for WSI analysis tasks.
@@
Most recently, state space models (SSMs), have introduced new opportunities for efficient representation learning in large-scale visual data [12, 11]. Vision Mamba, [49, 30, 40], an SSM-based architecture, enables modeling of short- and long-range spatial dependencies with linear computational complexity, all while maintaining favorable computational and memory characteristics, making it well-suited for WSI analysis. In this context, the application of modern, computationally efficient architectures to digital pathology represents a critical step toward advancing cancer prevention through improved risk stratification of precursor lesions.
@@
Conventional risk stratification methods categorize adenomas by size, histologic subtype, and grade of dysplasia, potentially overlooking finer-scale tissue patterns that may reflect underlying biological behavior. TAs with low-grade dysplasia are often regarded as biologically homogeneous, despite evidence that not all precancerous lesions progress to malignant disease. Subtle morphologic cues and variations, such as nuclear architecture, glandular organization, and epithelial maturation, may be difficult for human observers to consistently quantify but can be captured by deep learning models. This paper proposes XtraLight-MedMamba, an ultra-lightweight architecture incorporating ConvNext blocks, parallel Mamba-based layers, Spatial and Channel Attention Bridge (SCAB) and Fixed Non-Negative Orthogonal Classifier (FNOClassifier). ConvNeXt blocks extract shallow features from the input image and Mamba layers enhances and identifies subtle morphologic pattern cues through the modeling of both short- and long-range spatial dependencies. Then, SCAB module refines the intermediate feature representations. Whereas, FNOClassifier is a parameter-efficient classification module that classifies neoplastic TA features into “case” and “control” cohorts. The task-specific integration of these components into a compact architecture enables strong representation of both local glandular morphology and broader contextual patterns while maintaining an exceptionally low parameter count. Experimental results demonstrate that XtraLight-MedMamba outperforms previously established models, as detailed in subsequent sections.
@@
  • A Mamba-based ultra-lightweight architecture for classification of neoplastic tubular adenomas: XtraLight-MedMamba, combining the strengths of ConvNeXt blocks, mamba modules in parallel layers, a spatial and channel attention bridge module, and a fixed non-negative orthogonal classifier within a unified framework.
@@
  • A unique deep learning architectural strategy integrating (1) ConvNeXt blocks as a shallow feature extractor, acting upon design principles of Vision Transformers [dosovitskiy_image_2021] (ViT) such as depthwise separable convolutions, Layer Normalization, and MLP-style channel expansion to efficiently capture local morphologic patterns while preserving the spatial inductive bias of CNNs. (2) Parallel Vision Mamba (PVM) layers are included to model both short- and long-range spatial dependencies in parallel by branching out the features for state-space sequence modeling, each with channel $C/4$, enabling efficient global contextualization of morphologic features with linear computational complexity. (3) Multiscale feature fusion is enhanced by integration of SCAB modules by emphasizing relevant informative spatial regions and channel-wise responses. (4) FNOClassifier is incorporated to reduce redundant parameters deeper in the layers of the model while improving generalization and training stability through structured, orthogonal feature-to-class projections. This task-specific integration of these components into a compact architecture is tailored for subtle histopathologic pattern recognition.
  • Evaluation of the risk stratification performance of XtraLight-MedMamba on our newly introduced Neoplastic Tubular Adenoma (NPTA) dataset, a curated case-control dataset derived from WSIs of low-grade TAs.
@@
The proposed framework achieves strong classification performance with very low model complexity, yielding a favorable performance efficiency trade-off. By focusing on morphologic features critical for tissue characterization, XtraLight-MedMamba aims to improve the accuracy of neoplastic tubular adenoma classification by capturing subtle cellular and complex architectural patterns from case and control cohorts.
2 Related Work
Computational analysis of histopathology images has become a key research focus for improving CRC diagnosis and evaluation, driven by advances in computer vision. Early approaches relied on handcrafted algorithms that focused on specific image features to distinguish tissue types. Hamilton et al. [hamilton_automated_1997] introduced image texture analysis using co-occurrence matrix features and low optical density feature counts for classifying normal versus dysplastic tissues. This method automated the localization of dysplastic fields in colorectal histology. Subsequently, Kalkan et al. [kalkan_automated_2012] introduced an automated diagnosis of colorectal cancer from a whole biopsy slice that combined textural and structural features, using a two-level classification scheme. First, individual patches were classified as adenomatous, inflamed, cancerous, or normal using a k-nearest neighbors classifier. Afterwards, the slice-level information obtained from the patch distribution was then used to classify the slices using a logistic linear classifier. Another method proposed fully automated CRC grading from histology images [sengar_grading_2016]. The glands were first segmented automatically using an intensity-based thresholding approach, and the images were then classified as benign healthy, benign adenomatous, moderately differentiated malignant, or poorly differentiated malignant with a support vector machine classifier. While these methods were moderately effective for their respective tasks, they did not achieve the accuracy required to be viable tools for pathologists to detect and grade colorectal cancer from WSIs.
2.1 Deep Learning for Colorectal Cancer Histopathology
With the advent of deep learning, classical approaches have been largely replaced for automated CRC classification from histopathology images. CNNs were among the first deep learning architectures widely applied to histopathological image analysis, and they have been reviewed extensively alongside classical machine learning and CNN-based methods for classification, detection, and segmentation of relevant tissue [zhou_comprehensive_2020]. Gastrointestinal-related reviews further explored applications of deep learning for CRC detection and prognosis from histology slides [kuntz_gastrointestinal_2021]. Consequently, deep learning has become a major application in CRC histopathology, as evidenced by a systematic review of deep-learning-based CRC diagnosis [davri_deep_2022]. A transfer learning framework for classifying CRC histology images with sparse WSI annotations was proposed by Ben Hamida et al. [ben_hamida_deep_2021] and evaluated using a benchmark of multiple pre-trained CNN models. In their experiments, patch-level classification on their AiCOLO dataset was performed with an 18-layer ResNet [he_deep_2016] model. In addition, a multi-step segmentation model (UNet [ronneberger_u-net_2015], and SegNet [3]) was employed to generate semantic maps to identify abnormal regions on the slides. Meanwhile, Abdulrahman et al. [1] proposed an ensemble model combining EfficientNetV2 and DenseNet for binary classification of malignant versus benign colorectal tissue from WSIs using a custom Bahrain hospital dataset.
Many models are trained on fixed-size patches of WSIs, where the patch-level predictions can be aggregated to obtain slide-level classification [wang_accurate_2021, steimetz_deep_2025]. This strategy is computationally tractable for gigapixel-sized WSIs, applicable to a wide range of architectures, and has achieved strong performance across tasks, including cancer detection and tumor grading. For example, Wang et al. [wang_accurate_2021] used a weakly supervised, transfer-learned Inception-v3 for CRC detection and achieved slide-level classification accuracy on par with that of human pathologists. Paing and Pintavirooj [paing_adenoma_2023] proposed a ResNet-based architecture that applies a Fast Fourier Transform to a ResNet50 backbone for tumor grading, achieving similarly high performance. This model used cross-feature fusion to fuse local-scale spatial convolutions with global-scale Fourier convolutions. Another study by Steimetz et al. [steimetz_deep_2025] used an ImageNet-pretrained ResNet-34 model to classify between low-grade and high-grade dysplasia. Likewise, Zhou et al. [zhou_construction_2024] introduced a deep learning-based tumor risk signature (TRS) approach, using an ensemble of VGG19, ResNet50, and DenseNet21 models for the detection of stage III colorectal cancer. The model first segmented nine tissue types in WSIs, and subsequently, the tumor features were extracted to fit a Cox proportional hazards model. Similarly, DenseNetV2 [dragomir_quantitative_2025] was fine-tuned using the HALO image analysis platform to automatically detect the six morphotypes. The trained model was then applied to 644 sections from 161 cases to quantify morphotype areas. Furthermore, A lightweight, non-pretrained CNN [li_lightweight_2025] was proposed for the detection and visualization of multiclass colon tissue. It was integrated with a parametric Gaussian-distribution-based data-cleaning strategy to remove outliers and improve data quality. Another study employed a few-shot learning approach [li_few-shot_2024] that combined transfer learning and contrastive learning to classify CRC histopathology into benign and malignant categories. The model comprised modules for feature extraction, dimensionality reduction, and classification, and was trained using contrastive and cross-entropy losses. Subsequently, an ensemble deep learning model that combined the watershed algorithm to enhance glandular segmentation in colon histopathology was proposed by Roy et al. [roy_revolutionizing_2024]. This approach employed a UNet-based CNN, a weighted ensemble network that integrated DenseNet169 via augmentation, InceptionV3, and EfficientB3 as the backbone. Despite their success, conventional CNNs are inherently limited by the locality of convolutional operations.
To tackle this issue, Transformer-based architectures, enabled by the introduction of the Vision Transformer (ViT) [dosovitskiy_image_2021], have been explored to better capture long-range spatial relationships within colorectal tissue [sathyanarayana_colovit_2025, tp_deepcpd_2024, qin_colorectal_2024], and for the detection of higher-order structures within CRC [chen_predicting_2024]. Transformer-based architectures relied heavily on self-attention to capture such subtle patterns, which exhibit quadratic scaling [12, keles_computational_2023], making them difficult to scale efficiently.
ViT [dosovitskiy_image_2021] converted the input image into a set of patch-level embeddings, flattened them, and processed them like a sequence of tokens similar to Natural Language Processing (NLP) [43]. Using multi-head self-attention, the patches were then processed to capture information from across the image. While effective, ViT typically benefited from large datasets and high computational resources. To make transformers more scalable for images, Swin Transformer [liu_swin_2021], also known as Shifted Window Transformer was introduced, as an efficient variant of Transformers. Swin Transformers employed a hierarchical design, where attention was computed using non-overlapping local windows to minimize computational overhead. To maintain efficiency and prevent leakage of information, the window partitions were shifted between layers, allowing information to flow across neighboring windows.
In colonoscopy imaging, ColoViT [sathyanarayana_colovit_2025] was introduced as a hybrid of EfficientNet and ViT for advanced colon cancer detection. Here, EfficientNet extracted local features, whereas ViT captured global contextual information in colonoscopic images. Meanwhile, an efficient deep learning method, DeepCPD [tp_deepcpd_2024], was introduced to classify colonoscopic images into polyp versus non-polyp and hyperplastic versus adenoma. The model was a combination of a transformer and a Linear Multihead self-attention (LMSA) mechanism with data augmentation. In contrast, for histology, a hybrid of Swin Transformer for feature extraction and a skip-feedback connection with UNet was proposed to improve the model’s multi-level feature extraction capabilities, enabling end-to-end recognition of colorectal adenocarcinoma tissue images [qin_colorectal_2024]. Subsequently, Chen et al. [chen_predicting_2024] proposed DiNAT-MSI, an algorithm that incorporated the Dilated Neighborhood Attention Transformer (DiNAT) to enhance global context recognition and expand receptive fields, while avoiding fully quadratic global attention. Likewise, Transformer-Based Self-Supervised Learning and Distillation [li_transformer-based_2024] was introduced to improve Swin Transformer V2’s performance on CRC histology image classification, first, performing self-supervised pretraining and then fine-tuning with a layer-wise distillation technique on the NCT-CRC-HE-100K dataset. Furthermore, Guo et al. [guo_predicting_2023] proposed a Swin Transformer workflow to identify CRC molecular biomarkers directly from WSIs. It was also used to predict microsatellite instability from a small dataset.
2.2 Mamba for Cancer Classification
Subsequently, SSM-based architectures [12] like Vision Mamba [49] have been used to great effect in digital pathology tasks like WSI image classification [zhang_2dmamba_2025], due to their ability to capture global and local spatial context with far less computational power than transformers [zheng_m3amba_2025, ding_combining_2025]. 2DMamba [zhang_2dmamba_2025] based on 2D selective SSM framework was introduced by integrating the 2D spatial structure of images into Mamba. Coupled with a hardware-aware optimized operator, it preserved spatial continuity and delivered strong computational efficiency. Furthermore, a memory-driven Mamba network, M3amba [zheng_m3amba_2025], was introduced to address vanilla Mamba’s contextual forgetting when modeling long-range dependencies across thousands of instances. It consisted of a dynamic memory bank (DMB) that iteratively updated historical information, along with an intra-group bidirectional Mamba (BiMamba) block to improve feature representation and to fuse relevant historical information across groups, thereby facilitating richer inter-group connections. Another recent approach introduced a hybrid message-passing graph neural network (GNN) for local neighborhood interactions with Mamba to capture global tissue spatial relationships in WSIs [ding_combining_2025]. It was validated for predicting progression-free survival in patients with early-stage lung adenocarcinoma (LUAD). Moreover, SlideMamba [khan_slidemamba_2026], an entropy-based adaptive fusion of GNNs and Mamba, was introduced to enhance representation learning in digital pathology. It provided a principled mechanism for combining complementary feature streams and improving multi-scale representation learning by emphasizing the branch with lower predictive entropy. Next, Vim4Path [nasiri-sarvi_vim4path_2024], a self-supervised vision mamba, was introduced for evaluation on the Camelyon16 dataset for both patch-level and slide-level classification. The model used a Vision Mamba architecture, inspired by state-space models in the DINO framework, for representation learning.
While this line of SSM-based research has shown promising results in histopathology and digital pathology, there are few investigations on neoplasias of the colon, and even fewer on the prophylaxis of colorectal carcinoma. Existing studies have largely focused on well-studied disease settings, such as lung and breast cancer. Mamba-based modeling for risk stratification of precancerous lesions (adenomatous polyps), such as TAs, remains an underexplored diagnostic need. To bridge this gap, this work investigates an ultra-lightweight mamba-based architecture for the classification of neoplastic tubular adenoma in CRC WSI data. It aims to leverage the efficiency of SSMs and their capacity for short- and long-range contextual modeling while concurrently capturing diagnostically relevant glandular and stromal patterns.
2.3 Colorectal Cancer Histopathology Datasets
High-quality datasets have played a central role in advancing automated CRC classification methods from histopathology images. Most datasets consist of hematoxylin and eosin (H&E)-stained WSIs, considered the “gold standard” for histopathology, that are divided into smaller patches and annotated by expert pathologists with tissue-type or prognostic labels at either the patch or slide level. While many studies use institution-specific, non-public datasets [davri_deep_2022], several publicly available datasets have been used for training and evaluation of CRC-related vision tasks. The NCT-CRC-HE-100K dataset comprises 100,000 labeled image patches from CRC tissue slides spanning nine colorectal tissue classes [kather_100000_2018] and is widely used for supervised training. For tumor grading, Barbano et al. [barbano_unitopatho_2021] introduced the UniToPatho dataset in 2021, consisting of nearly 10,000 image patches from six different classes corresponding to tissue type and dysplasia grade, including low- and high-grade dysplasia for TAs.
In contrast to these datasets, our Neoplastic Tubular Adenoma (NPTA) dataset [39], [31] is designed as a retrospective cohort consisting solely of low-grade dysplastic TAs identified at screening colonoscopy. The dataset is released as an expertly annotated set of image patches derived from high-resolution WSIs of TA, stratified into two groups based on the outcomes of their follow-up screening(s): case and control. Case slides are derived from patients who subsequently developed CRC, and control slides are derived from patients who did not. This case–control design enables deep learning models such as XtraLight-MedMamba to learn subtle, early morphologic signals, perhaps associated with future CRC risk, from lesions that are not necessarily considered “high-risk” by current routine histologic guidelines.